OpenStudy (ammarah):

Solve the system using any algebraic method!!!!!! HELP I BEG U 2x+y-z=-1 x+2y+2z=10 2x+z=3

hero (hero):

Use matrices

OpenStudy (ammarah):

no i have to use elimination can u show me step by step how do do it please ?

hero (hero):

I thought it said to use any algebraic method...but now you have to use elimination method only?

OpenStudy (ammarah):

yeah our teacher told us to use elimination......can u please draw out the stepss?

OpenStudy (anonymous):

I'll do it haha

OpenStudy (anonymous):

We have: \[2x+y-z=-1\phantom{..}\rightarrow\phantom{..}y=z-1-2x\] We also have: \[\eqalign{ &x+2y+2z=10 \\ &x+2(z-1-2x)=10 \\ &x+2z-2-4x=10 \\ &-3x+2z=12 \\ &2z=12+3x \\ &z=\frac{1}{2}(12+3x) \\ }\] We cal also redefine y as a result of isolating for \(z\): \[\eqalign{ y&=z-1-2x \\ &=(1/2)(12+3x)-1-2x \\ &=6+(3/2)x-1-2x\\ &=5-(1/2)x }\] So now we have z and y in terms of x, let's slove for x: \[\eqalign{ &2x+z=3 \\ &2x+(1/2)(12+3x)=3 \\ &2x+6+(3/2)x=3 \\ &(7/2)x=-3 \\ &x=-\frac{6}{7} \\ }\] So now that we have x, and since y and z were in terms of x, we can find z and y: \[z=\frac{1}{2}(12+3x)=\frac{1}{2}\left(12+3*-\frac{6}{7}\right)=\frac{1}{2}\left(12-\frac{18}{7}\right)=\frac{1}{2}\times\frac{66}{7}=\frac{66}{14}\] \[y=5-\frac{1}{2}x=5-\frac{1}{2}\times-\frac{6}{7}=5+\frac{6}{14}=\frac{76}{14}\]

OpenStudy (anonymous):

So therefore, the solution to the system \(S\): \[S=\left\{\eqalign{ &2x+y-z=-1 \\ &x+2y+2z=10\\ &2z+x=3\\ }\right\}\] Is the single point: \[SOLUTION_{\{S\}}=\left(x,y,z\right)=\left(-\frac{6}{7},\frac{76}{14},\frac{66}{14}\right)\]

OpenStudy (anonymous):

WHEW! Cool?